OpenStudy (anonymous):

5x + 8y=-38 4x + 7y=-34 (substitution)

OpenStudy (anonymous):

Use on of the equations to get x in respect to y I'll use the first one: 5x + 8y = -38 5x = -8y - 38 x = (-8y - 38)/5 Substitute this in the second equation, let's see what you get

OpenStudy (anonymous):

I'm no good at this.... :(

OpenStudy (anonymous):

You've got to try at least. If you make a wrong, we can together make it right. Just sitting down there and letting me do all the work wouldn't be nice, would it?

OpenStudy (anonymous):

4x+7(3x+12)=34 4x+21x+84=34 25x=34-84 25x=-50 x=-50/25 x=-2 answer. 4*-2+7y=34

OpenStudy (anonymous):

his answer looks different from your example...why is that

OpenStudy (anonymous):

shereenkhan , do you mind explaining what you did ? I'd very much be interested. How did you get here --> 7(3x+12) ?

OpenStudy (anonymous):

marucs13 , let's see what he's got for us. I have no idea, the angle he's pitching.

OpenStudy (anonymous):

yeah...i'm so confused...

OpenStudy (anonymous):

See, he just cowed out! Let's continue from where we were.

OpenStudy (anonymous):

This is our derived eqn for x x = (-8y - 38)/5 We'll subst. that value in the second eqn 4x + 7y=-34 4(-8y - 38)/5 + 7y=-34

OpenStudy (anonymous):

Are you following?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now, we're going to multiply through by 5 to remove that fraction bit there 4(-8y - 38)/5 * 5 + 7y * 5=-34 * 5 yields: 4(-8y - 38) + 35y = -170 expanding the bracket: -32y - 152 + 35y = -170 --> 35y - 32y = -170 + 152 3y = -18 div both sides by 3 y = -6

OpenStudy (anonymous):

ah i get it :)

OpenStudy (anonymous):

So, you know how to get x now?

OpenStudy (anonymous):

oh..no

OpenStudy (anonymous):

Remember our derived eqn for x? x = (-8y - 38)/5 subst. y = -6 into the eqn and let's see what you got.

OpenStudy (anonymous):

ah i thought so...i was thinking i had to substitute y=-6 in but i didn't know where... so x= (-6)/5?

OpenStudy (anonymous):

No, where you have y in the eqn, just put -6 instead

OpenStudy (anonymous):

x = (-8(-6) - 38)/5?

OpenStudy (anonymous):

good! evaluate that and see what x is

OpenStudy (anonymous):

x=(48 - 38)/5 x=(10)/5 x=2

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

really?!? :D

OpenStudy (anonymous):

So, x = 2, y = -6

OpenStudy (anonymous):

so, those are the two answers i use?

OpenStudy (anonymous):

Of course!

OpenStudy (anonymous):

but...how would i show my work...there's so much here

OpenStudy (anonymous):

Just go through every step we've worked so far, and follow them logically... good luck! :)